### v1

# First-Principles Framework for the Prediction of Intersystem Crossing Rates in Spin Defects: The Role of Electron Correlation

Abstract
Optically active spin defects in solids are promising platforms for quantum technologies. Here, we present a first-principles framework to investigate intersystem crossing processes, which represent crucial steps in the optical spin-polarization cycle used to address spin defects. Considering the nitrogen-vacancy center in diamond as a case study, we demonstrate that our framework effectively captures electron correlation effects in the calculation of many-body electronic states and their spin-orbit coupling and electron-phonon interactions, while systematically addressing finite-size effects. We validate our predictions by carrying out measurements of fluorescence lifetimes, finding excellent agreement between theory and experiments. The framework presented here provides a versatile and robust tool for exploring the optical cycle of varied spin defects entirely from first principles.
Introduction—Optically active spin defects in solids offer a promising platform for advancing quantum technologies [1]. Prototypical systems, such as the negatively charged nitrogen-vacancy (NV-) center in diamond [2, 3] and neutral divacancy centers in silicon carbide [4], have shown potential applications in quantum sensing [5, 6], communication [7, 8, 9, 10], and computation [11, 12] by functioning as spin qubits. First-principles calculations have been instrumental in studying these spin defects [13, 14, 15, 16], in particular their optical spin-polarization cycle, which is critical for the initialization and readout of qubit states [17, 18, 19]. While radiative transitions have been extensively studied using first-principles approaches [20, 21, 22], non-radiative intersystem crossing (ISC) transitions remain less explored. A central challenge lies in accurately accounting for electron correlation effects, which are important for the description of many-body electronic states and their associated spin-orbit coupling (SOC) and electron-phonon (e-ph) interactions [23].
In this Letter, we present a theoretical and computational framework to predict ISC rates in spin defects, highlighting the key role of electron correlation in the accurate description of the defects’ many-body states, their SOC, and e-ph interactions. We focus on the NV- center in diamond as an exemplar system and present results for ISC rates as a function of temperature. We validate our theoretical predictions by measuring fluorescence lifetimes, achieving excellent agreement between theory and experiment.
The NV- center, comprising of a nitrogen substitution and an adjacent vacancy, exhibits three localized defect orbitals ($a_{1}$, $e_{x}$, $e_{y}$) within the diamond band gap. In the negatively charged state, four electrons occupy these defect orbitals, resulting in a triplet ground state (${}^{3}\!A_{2}$), a triplet excited state (${}^{3}\!E$), and two singlet states (${}^{1}\!E$ and ${}^{1}\!A_{1}$[24, 25, 26]. The optical spin-polarization cycle (Fig. 1) consists of excitation from the triplet ground to the triplet excited state, followed by ISC to the higher singlet state, with relaxation to a lower singlet state, and finally an ISC process back to the triplet ground state. This cycle enables the initialization and readout of the qubit states, which are defined by the spin sublevels of the ${}^{3}\!A_{2}$ ground state [18].
First-principles calculations have been successfully employed to investigate the optical transitions between ${}^{3}\!A_{2}$ and ${}^{3}\!E$ [20, 27, 21] and between ${}^{1}\!A_{1}$ and ${}^{1}\!E$ [18, 22]. However, studies of ISC processes remain limited due to the challenges of accurately capturing electron correlation while also addressing finite-size effects. Previous ISC studies relying on mean-field approaches, e.g., density functional theory (DFT), overestimate measured ISC rates due to an inadequate description of SOC and e-ph interactions [17, 18]. On the other hand, advanced quantum chemistry methods, applied to hydrogen-terminated carbon clusters simulating an NV- in diamond, underestimate ISC rates. The underestimate likely arises from quantum confinement effects and the negative electron affinity of small carbon clusters, which lead to inaccurate descriptions of electronic states and SOC [28, 29, 30]. Moreover, cluster models are unsuitable for the direct simulation of e-ph interactions in the bulk environment.
Here, we focus on ISC transitions from the $\widetilde{A}_{1}$ and $\widetilde{E}_{1,2}$ vibronic levels of ${}^{3}\!E$ to ${}^{1}\!A_{1}$ ($\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ in Fig. 1) and demonstrate the capabilities of our general theoretical and computational framework fully based on first-principles calculations. We combine several methods recently developed to investigate electronic structure properties beyond mean-field approaches. Specifically, we compute SOC using many-body wavefunctions from the quantum defect embedding theory (QDET) [31, 32, 33, 34], which accounts for electron correlation within the active space formed by defect states, while incorporating environmental effects through the screened Coulomb interaction. E-ph interactions are evaluated by computing the vibrational overlap function (VOF) between the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states, using atomic geometries and phonon modes obtained from spin-conserving and spin-flip time-dependent DFT (TDDFT) calculations [35]. This method effectively accounts for the multi-configurational nature of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and has been previously shown to accurately predict the absorption spectrum between singlet states of the NV- center [22]. Additionally, the dynamic Jahn-Teller (DJT) [36] and Herzberg-Teller (HT) effects [37, 38] are incorporated into the VOF calculations. Importantly, our framework relies exclusively on solid-state calculations and is scalable to systems containing hundreds to thousands of atoms, thereby enabling a systematic treatment of finite-size effects for all computed quantities. We first describe the various steps of the computational protocol, and then we compare fluorescence lifetimes with measured values.
Figure 1: Many-body states and optical spin polarization cycle of the NV- center in diamond. The spin-orbit splitting $\lambda_{z}$ represents the energy separation between the sublevels of the ${}^{3}\!E$ state. The intersystem crossing (ISC) rates from vibronic levels $\widetilde{A}_{1}$ and $\widetilde{E}_{1,2}$ to ${}^{1}\!A_{1}$ are labeled as $\Gamma_{A_{1}}$ (red dashed line) and $\Gamma_{E_{1,2}}$ (yellow dashed line). Radiative transitions ${}^{3}\!E\to{}^{3}\!A_{2}$ (blue solid line) and ${}^{1}\!A_{1}\to{}^{1}\!E$ (marron solid line), non-radiative transition ${}^{1}\!A_{1}\to{}^{1}\!E$ (marron dashed line), and ISC ${}^{1}\!E\to{}^{3}\!A_{2}$ (purple dashed line) are shown for completeness.
Theoretical framework—At temperatures near 0 K, the ISC rate $\Gamma_{A_{1}}$ is calculated using Fermi’s golden rule:
$\Gamma_{A_{1}}=\dfrac{4\pi}{\hbar}|\lambda_{\perp}|^{2}F_{A_{1}}\left(\Delta\right),$ (1)
where $\lambda_{\perp}=1/\sqrt{2}\left\langle{}^{3}\!E(A_{1})\right|\hat{H}_{\mathrm{SO}}\left|{}^{1}\!A_{1}\right\rangle$ is the SOC matrix element between the $A_{1}$ electronic sublevel of ${}^{3}\!E$ and ${}^{1}\!A_{1}$ [25, 26]. Another matrix element, $\lambda_{z}=\left\langle{}^{3}\!E(A_{1})\right|\hat{H}_{\mathrm{SO}}\left|{}^{3}\!E(A_{1})\right\rangle$, determines the splitting of the sublevels in ${}^{3}\!E$. The term $F_{A_{1}}(\Delta)$ is the VOF for the transition from the $\widetilde{A}_{1}$ vibronic level of ${}^{3}\!E$ to the various vibronic levels of ${}^{1}\!A_{1}$. The energy $\Delta$ represents the gap between $\widetilde{A}_{1}$ and the lowest vibronic level of ${}^{1}\!A_{1}$. A detailed derivation of Eq. 1 is provided in Sec. S5 [39]. The ISC rate $\Gamma_{E_{1,2}}$ can be similarly obtained by using the VOF for the $\widetilde{E}_{1,2}$ vibronic level [$F_{E_{1,2}}(\Delta)$].
To properly include electron correlation effects, $\lambda_{z/\perp}$ are computed using many-body wavefunctions obtained from QDET calculations [31, 32, 33, 34] and using the many-body SOC operator, defined as [40, 41]
$\hat{H}_{\mathrm{SO}}=\sum_{i=1}^{N_{e,\mathcal{A}}}\hat{\mathbf{z}}_{i}\cdot\hat{\mathbf{s}}_{i}=\sum_{pq}\sum_{\sigma\sigma^{\prime}}\langle p|\hat{\mathbf{z}}|q\rangle\langle\sigma|\hat{\mathbf{s}}\left|\sigma^{\prime}\right\rangle\hat{c}_{p\sigma}^{\dagger}\hat{c}_{q\sigma^{\prime}},$ (2)
where $N_{e,\mathcal{A}}$ is the number of electrons in the QDET active space, and $\hat{c}^{\dagger}_{p\sigma}$ ($\hat{c}_{q\sigma^{\prime}}$) are the creation (annihilation) operators for orbital $p$ ($q$) in spin channel $\sigma$ ($\sigma^{\prime}$). The one-body SOC operators entering Eq. 2 are defined as
$\hat{h}_{\mathrm{SO}}=\hat{\mathbf{z}}\cdot\hat{\mathbf{s}}=\sum_{I=1}^{N_{\mathrm{atom}}}\left(\hat{V}_{\mathrm{NL},I}^{\mathrm{FR}}-\hat{V}_{\mathrm{NL},I}^{\mathrm{SR}}\right),$ (3)
where $\hat{V}_{\mathrm{NL},I}^{\mathrm{FR}}$ and $\hat{V}_{\mathrm{NL},I}^{\mathrm{SR}}$ are the non-local (NL) components of the fully relativistic (FR) and the scalar relativistic (SR) pseudopotentials, respectively, for atom $I$ in the supercell [42, 43, 44]. Finite-size effects are addressed by converging the values of $\lambda_{z/\perp}$ with supercell size.
Figure 2: Computed spin-orbit coupling (SOC) parameters $\lambda_{z/\perp}$. (a) $\lambda_{z/\perp}^{\mathrm{DFT}}$ as a function of supercell size, with the extrapolated values ($\lambda_{z/\perp,0}^{\mathrm{DFT}}$) shown as dashed lines. $\lambda_{z/\perp}^{\mathrm{DFT}}$ is obtained using group-theory-derived many-body wavefunctions with a minimal defect orbital model ($a_{1},e_{x},e_{y}$). (b) $\lambda_{z/\perp}^{\mathrm{QDET}}$, computed using the many-body wavefunctions from QDET calculations in a 511-atom supercell with varying active spaces. The shaded areas indicate uncertainties due to active space selection. (c) Comparison of SOC parameters from this work (DFT and QDET) with previous theoretical studies [17, 29, 45] and estimated experimental values [46, 47, 48, 49, 25]. Full data available in Table S3 [39].
The VOF $F_{A_{1}/E_{1,2}}(\Delta)$ is computed in two steps. First, contributions from $a_{1}$-type phonon modes, $F^{a_{1}}(\Delta)$, are calculated using the Huang-Rhys (HR) theory [50, 20, 21]. This calculation uses the atomic geometries of ${}^{3}\!E$ and ${}^{1}\!A_{1}$ and the phonon modes of ${}^{1}\!A_{1}$, derived from spin-conserving and spin-flip TDDFT calculations [22], which effectively include the multi-configurational nature of these states. Next, the DJT effect of ${}^{3}\!E$, involving $e$-type phonon modes, is incorporated to obtain the final $F_{A_{1}/E_{1,2}}(\Delta)$ [17]. The computed VOFs are extrapolated to the dilute limit using the force constant embedding approach [20, 27]. Equation 1 assumes the validity of the Franck-Condon (FC) principle and treats $\lambda_{\perp}$ as independent of atomic vibrations. Here, we remove this assumption by accounting for the HT effect [37, 38] and compute the derivative of $\lambda_{\perp}$ with respect to configuration coordinates in the calculation of $F^{a_{1}}(\Delta)$.
The energy gap $\Delta$ is approximated as the adiabatic energy difference between ${}^{3}\!E$ and ${}^{1}\!A_{1}$ at their respective equilibrium atomic geometries, assuming approximate cancellation of their zero-point energies. The vertical energy gap between ${}^{3}\!E$ and ${}^{1}\!A_{1}$ is computed using QDET, while their FC shifts (reorganization energies) are obtained from TDDFT [22, 35]. Further details on our computational framework are in Secs. S1 and S2 [39].
Results—We first compute the SOC parameter, $\lambda_{z/\perp}$, using many-body wavefunctions derived from group theory, using the minimal basis set of defect orbitals $a_{1}$, $e_{x}$, and $e_{y}$ (see Tables S1 and S2 [39]). The resulting expressions are: $\lambda_{z}=|\langle e_{x}|\hat{\mathbf{z}}|e_{y}\rangle|/2$ and $\lambda_{\perp}=|\langle a_{1}|\hat{\mathbf{z}}|e_{x}\rangle|/2$. Results obtained from these expressions are denoted as $\lambda_{z/\perp}^{\mathrm{DFT}}$. Figure 2(a) shows the strong dependence of $\lambda_{z/\perp}^{\mathrm{DFT}}$ on supercell size, attributed to the overlap of defect orbitals with periodic images in small supercells. Extrapolating to the dilute limit via exponential decay yields $\lambda_{z,0}^{\mathrm{DFT}}=37.5$ GHz and $\lambda_{\perp,0}^{\mathrm{DFT}}=70.6$ GHz.
To properly account for electron correlation effects beyond the group theory analysis, we compute $\lambda_{z/\perp}$ using many-body wavefunctions from QDET calculations, denoted as $\lambda_{z/\perp}^{\mathrm{QDET}}$. Figure 2(b) presents $\lambda_{z/\perp}^{\mathrm{QDET}}$ for a 511-atom supercell, highlighting the significant influence of electron correlation. While QDET wavefunctions resemble those from the group theory analysis, they do contain additional configurations that introduce small yet important corrections, leading to a notable decrease in $\lambda_{z/\perp}^{\mathrm{QDET}}$ compared to $\lambda_{z/\perp}^{\mathrm{DFT}}$ (details in Secs. S3 and S4 [39]). By accounting for the weak dependence of $\lambda_{z/\perp}^{\mathrm{QDET}}$ on the active space choice and applying a finite-size correction (assumed to be the same as for $\lambda_{z/\perp}^{\mathrm{DFT}}$), we determine the final SOC parameters as $\lambda_{z,0}^{\mathrm{QDET}}=23.1\pm 4.3$ GHz and $\lambda_{\perp,0}^{\mathrm{QDET}}=29.4\pm 2.8$ GHz. These values are used in the ISC rate calculations presented below.
In Fig. 2(c), we compare our computed $\lambda_{z/\perp}$ with previous theoretical and experimental results. The quantity $p\lambda_{z}$ represents the measured fine structure splitting of ${}^{3}\!E$, where $p$ accounts for the Ham reduction factor due to the DJT effect. Using $p=0.304$ from Ref. [17], we obtain $p\lambda_{z,0}^{\mathrm{QDET}}=7.0\pm 1.3$ GHz, in good agreement with the measured values: $5.33\pm 0.03$ GHz [46, 47], whereas $p\lambda_{z,0}^{\mathrm{DFT}}$ overestimates experiment by a factor of two. Our computed $\lambda_{\perp,0}^{\mathrm{QDET}}$ also aligns well with the experimental estimate based on $\lambda_{\perp}/\lambda_{z}=1.2\pm 0.2$ relation [48, 49], while $\lambda_{\perp,0}^{\mathrm{DFT}}$ is approximately three times as large. We note that previous quantum chemistry CASSCF calculations [29] significantly underestimated $\lambda_{\perp}$, likely due to an inaccurate description of the NV- center’s many-body wavefunctions with a small C33H36N- cluster.
As mentioned above, in our calculations of VOFs for the $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ ISC transitions, we first considered contributions from $a_{1}$-type phonon modes. The computed $F^{a_{1}}(\Delta)$ is shown in Fig. 3(a). Previous studies often derived $F^{a_{1}}(\Delta)$ from constrained-occupations DFT ($\Delta$SCF), which approximates ${}^{3}\!E$ and ${}^{1}\!A_{1}$ as single Slater-determinant states and inadequately account for electron correlation effects [17]; alternatively, $F^{a_{1}}(\Delta)$ was approximated by the VOF for the ${}^{3}\!E\to{}^{3}\!A_{2}$ photoluminescence (PL) transition [48, 49, 29]. Here, we compute $F^{a_{1}}(\Delta)$ using the HR theory [50], based on equilibrium atomic geometries of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and $a_{1}$-type phonon modes of ${}^{1}\!A_{1}$, obtained from spin-conserving and spin-flip TDDFT calculations. A scaling factor is applied to adjust the HR factors, correcting for underestimated atomic displacements due to the use of the PBE functional [27, 22] (see Sec. S6 [39]). We find notable differences between $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{1}\!A_{1}$ and ${}^{3}\!E\to{}^{3}\!A_{2}$ transitions [Fig. 3(a)]. Specifically, in the range $\Delta\in[0.3,0.4]$ eV, $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{3}\!A_{2}$ exceeds that of ${}^{3}\!E\to{}^{1}\!A_{1}$ by a factor of 1.5–1.9. Furthermore, the ${}^{3}\!E\to{}^{3}\!A_{2}$ PL lineshape overestimates $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{1}\!A_{1}$ by a factor of 1.9–3.1 due to contributions from $e$-type phonon modes.
Figure 3: Vibrational overlap functions (VOFs). (a) Computed VOFs for the ${}^{3}\!E\to{}^{1}\!A_{1}$ and ${}^{3}\!E\to{}^{3}\!A_{2}$ transitions under the Franck-Condon (FC) principle, considering $a_{1}$-type phonon modes and extrapolated to the dilute limit. The experimental photoluminescence (PL) spectral function for ${}^{3}\!E\to{}^{3}\!A_{2}$ [20] is shown as the gray shaded area. (b) VOFs for the ${}^{3}\!E\to{}^{1}\!A_{1}$ transition with $a_{1}$-type phonon contributions, showing FC (dashed), Franck-Condon Herzberg-Teller (FCHT, dashed-dotted), and Herzberg-Teller (HT, dotted) components, with the total function as the solid line. (c) VOFs for $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ intersystem crossing transitions, incorporating the dynamic Jahn-Teller effect of the ${}^{3}\!E$ state.
The VOF computed using the HR theory assumes the validity of the FC principle, where $\lambda_{\perp}$ is treated as constant between the atomic geometries of the initial and final states of the ISC transition. However, we find that $\lambda_{\perp}$ varies by 8% between the ${}^{1}\!A_{1}$ and ${}^{3}E$ equilibrium atomic geometries (see Sec. S7 [39]). This variation is accounted for by including the HT effect [37, 38] in our calculations. Using the ${}^{1}\!A_{1}$ equilibrium atomic geometry as the reference, we compute the Franck-Condon Herzberg-Teller (FCHT) and HT VOFs, shown in Fig. 3(b). Our results reveal that while the HT term contributes only 1%, the FCHT term reduces the FC term intensity by 19%. These findings indicate that the HT effect has a minor influence on the VOF in the NV- center. However, our calculations demonstrate the capability of the framework developed here to assess vibronic coupling effects in SOC calculations [51].
Finally, using an effective Hamiltonian describing the interaction between ${}^{3}\!E$ electronic states and $e$-type phonon modes [17], we incorporate the DJT effect into the VOF calculations for $\Gamma_{A_{1}}$ and the $\Gamma_{E_{1,2}}$, yielding $F_{A_{1}}(\Delta)$ and $F_{E_{1,2}}(\Delta)$ [Fig. 3(c)]. The DJT effect mixes the $A_{1}$, $A_{2}$, and $E_{1,2}$ electronic levels of ${}^{3}\!E$, giving rise to the vibronic states $\widetilde{A}_{1}$, $\widetilde{A}_{2}$, and $\widetilde{E}_{1,2}$ through coupling with $e$-type phonons. This coupling reduces the intensity of $F_{A_{1}}(\Delta)$ relative to the bare $F^{a_{1}}(\Delta)$ and generates a non-zero $F_{E_{1,2}}(\Delta)$, thereby enabling the $\Gamma_{E_{1,2}}$ ISC transition. The onset of $F_{E_{1,2}}(\Delta)$ at 78 meV confirm its $e$-type phonon-assisted nature, consistent with the model of Ref. [48]. Further details are provided in Sec. S5 [39].
We now combine the calculations of the SOC parameter $\lambda_{\perp,0}^{\mathrm{QDET}}$ and the VOFs $F_{A_{1}}(\Delta)$ and $F_{E_{1,2}}(\Delta)$ to compute the ISC rates $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ as functions of the energy gap $\Delta$. Results for $\Gamma_{A_{1}}$ are shown in Fig. 4(a), while those for $\Gamma_{E_{1,2}}$ are in Fig. S10 [39]. Comparing the computed $\Gamma_{A_{1}}$ with experimental data [52, 48] yields an estimated $\Delta$ range of $[0.334,0.389]$ eV. Similarly, the comparison of $\Gamma_{E_{1,2}}$ with experimental results [52, 48] gives $[0.314,0.375]$ eV. The overlap of these two ranges provides a refined estimate of $\Delta\in[0.334,0.375]$ eV, offering a reliable prediction for the ${}^{3}\!E-{}^{1}\!A_{1}$ energy gap. The ISC rates computed using $\lambda_{\perp,0}^{\mathrm{DFT}}$, also shown in Fig. 4, overestimate experiments by roughly a factor of 6. The inverse dependence of $\Gamma_{A_{1}/E_{1,2}}$ on $\Delta$ suggests that the NV- center’s ISC rates, and hence the optical spin-readout contrast, could be enhanced by decreasing $\Delta$, potentially through strain engineering [53, 54].
Our QDET calculations predict a $\Delta$ value of 0.414 eV, which exceeds the upper bound of $\Delta$ shown in Fig. 4(a), resulting in a $\Gamma_{A_{1}}$ rate of $66\pm 13$ MHz. This rate is slightly lower than the measured value of $100.5\pm 3.8$ MHz [48]. We also compare previously reported $\Delta$ values from various theoretical approaches [35, 55, 56, 57, 28, 29, 58] and compute the corresponding $\Gamma_{A_{1}}$ rates using $\lambda_{\perp,0}^{\textrm{QDET}}$ and $F_{A_{1}}(\Delta)$ from this work, as shown in Fig. 4(b). Notably, different theoretical approaches predict $\Delta$ in a broad range of $[0.04,0.51]$ eV, leading to $\Gamma_{A_{1}}$ between 20 MHz and 450 MHz. The values of $\Delta$ obtained from quantum chemistry approaches using carbon clusters are sensitive to both the level of theory and cluster size, resulting in $\Gamma_{A_{1}}$ variations by an order of magnitude. Overall, the comparison in Fig. 4(b) provides a valuable benchmark for quantitatively assessing the performance of various theoretical methods in predicting the energies of many-body states of spin defects. (See Sec. S9 [39] for a detailed comparison of ISC rates evaluated here and in previous studies [48, 49, 17, 29]).
Figure 4: Intersystem crossing (ISC) rates and fluorescence lifetimes. (a) Computed ISC rate $\Gamma_{A_{1}}$ as a function of the energy gap ($\Delta$) between ${}^{3}\!E$ and ${}^{1}\!A_{1}$, with uncertainty (red shaded area) from the SOC parameter $\lambda_{\perp,0}^{\text{QDET}}$. The bounds $\Delta_{-}$ and $\Delta_{+}$ are determined by the intersection of the computed $\Gamma_{A_{1}}$ with experimental results [52]. $\Gamma_{A_{1}}$ from $\lambda_{\perp,0}^{\text{DFT}}$ is shown for comparison (dotted line). (b) $\Gamma_{A_{1}}$ computed using $\lambda_{\perp,0}^{\text{QDET}}$ and the vibrational overlap function $F_{A_{1}}(\Delta)$ from this work, with $\Delta$ values from various theoretical methods [35, 55, 56, 57, 28, 29, 58], including Franck-Condon shifts from TDDFT [35]. A visualization width of 0.02 eV is applied to each $\Delta$. Detailed data are provided in Table S5 [39]. (c) Computed fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ compared with experimental values from this work and previous studies by Batalov et al. [59], Robledo et al. [60], and Toyli et al. [61].
Finally, we extend our computational framework for ISC rate calculations to finite temperatures ($T$) by summing contributions from different vibronic levels of ${}^{3}\!E$, weighted by the Boltzmann factor (see Sec. S8 [39]). We compute the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 K to 700 K, defined as $\tau(T)=1/\left[\Gamma_{\mathrm{Rad}}+\Gamma_{\mathrm{ISC}}(T)\right]$, where $\Gamma_{\mathrm{Rad}}=82.9$ MHz is the radiative decay rate of ${}^{3}\!E$ [52], and $\Gamma_{\mathrm{ISC}}(T)$ is the average ISC rate at a given $T$. To validate our calculations, we measured the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 K to 300 K in 50 K intervals. Combining these with previous measurements from Batalov et al. [59], Robledo et al. [60], and Toyli et al. [61], we obtain experimental data spanning 100–600 K. Details on the experimental setup and lifetime fitting are provided in Secs. S10 and S11 [39]. Despite moderate uncertainties in measurements and lifetime-fitting, our computed lifetimes agree well with experimental values below 600 K, as shown in Fig. 4(c). At temperatures above 600 K, a significant decrease in lifetime is observed experimentally, which may be attributed to additional decay pathways [61, 48] not yet included in our calculations, which will be explored in future work.
Discussion—In summary, we have developed a general theoretical and computational framework for the evaluation of ISC rates in spin defects, enabling a complete description of their optical cycle based on high-level first-principles calculations. Using the NV- center in diamond as a case study, we have demonstrated that our framework accurately accounts for electron correlation in many-body states, thus enabling precise calculations of SOC parameters and e-ph interactions. We have also incorporated DJT and HT effects in VOF calculations, yielding ISC rates in good agreement with experiments. We have validated our approach at finite temperatures by comparing computed fluorescence lifetimes with experimental data, achieving excellent agreement between theory and experiment. Importantly, unlike cluster model-based methods [28, 62, 63, 29], our framework is entirely based on calculations for bulk systems and is scalable to systems containing hundreds to thousands of atoms, facilitating a systematic evaluation of finite-size effects. We emphasize that the framework presented here is general and applicable to studying ISC mechanisms and optical spin-polarization cycles in broad classes of spin defects; hence, it offers a robust tool for interpreting experiments under external fields, and guiding the engineering of spin defects, e.g., through the refinement of high-throughput discovery workflows.
Promising directions for future work include extending the current approach to include a multi-phonon-mode treatment of DJT and HT effects for the study of vibronic SOC [51, 27, 64]. Additionally, our framework can be adapted to calculate phosphorescence rates for transition metal impurities in solids [65], further expanding the applicability of first-principles calculations in predicting spin-related phenomena in quantum materials.
Acknowledgments—The authors would like to thank Yuhang Ai for helpful discussions. The theoretical and computational work was primarily supported by the Midwest Integrated Center for Computational Materials (MICCoM) as part of the Computational Materials Sciences Program funded by the U.S. Department of Energy (Y.J., J.P., C.E. B.H., M.G., F.J.H., and G.G.). J.P. acknowledges the support from the Chicago Prize Postdoctoral Fellowship in Theoretical Quantum Science. Additional support for experimental validation includes the U.S. Department of Energy, Office of Science, Basic Energy Sciences Materials Sciences and Engineering Division (B.P. and D.D.A.), the SNU-Global Excellence Research Center establishment project and the NRF grant funded by the Korean government (D.D.O.) (MSIT) No. RS-2023-00258359, and from the AFOSR MURI under award No. FA9550-23-1-0330 (M.M.M. and C.B.). This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under contract No. DE-AC02-05CH11231 using NERSC Award No. ALCC-ERCAP0025950, and resources of the University of Chicago Research Computing Center.

### v2

@@
 First-principles calculations have been instrumental in studying these spin defects [13, 14, 15, 16], in particular their optical spin-polarization cycle, which is critical for the initialization and readout of qubit states [17, 18, 19].
-While radiative transitions have been extensively studied using first-principles approaches [20, 21, 22], non-radiative intersystem crossing (ISC) transitions remain less explored.
+While radiative transitions have been extensively studied using first-principles approaches [20, 21, 22], nonradiative intersystem crossing (ISC) transitions remain less explored.
 A central challenge lies in accurately accounting for electron correlation effects, which are important for the description of many-body electronic states and their associated spin-orbit coupling (SOC) and electron-phonon (e-ph) interactions [23].
@@
 We validate our theoretical predictions by measuring fluorescence lifetimes, achieving excellent agreement between theory and experiment.
-The NV- center, comprising of a nitrogen substitution and an adjacent vacancy, exhibits three localized defect orbitals ($a_{1}$, $e_{x}$, $e_{y}$) within the diamond band gap.
+The NV- center, comprising a nitrogen substitution and an adjacent vacancy, exhibits three localized defect orbitals ($a_{1}$, $e_{x}$, $e_{y}$) within the diamond band gap.
 In the negatively charged state, four electrons occupy these defect orbitals, resulting in a triplet ground state (${}^{3}\!A_{2}$), a triplet excited state (${}^{3}\!E$), and two singlet states (${}^{1}\!E$ and ${}^{1}\!A_{1}$[24, 25, 26].
@@
 E-ph interactions are evaluated by computing the vibrational overlap function (VOF) between the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states, using atomic geometries and phonon modes obtained from spin-conserving and spin-flip time-dependent DFT (TDDFT) calculations [35].
-This method effectively accounts for the multi-configurational nature of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and has been previously shown to accurately predict the absorption spectrum between singlet states of the NV- center [22].
+This method effectively accounts for the multiconfigurational nature of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and has been previously shown to accurately predict the absorption spectrum between singlet states of the NV- center [22].
 Additionally, the dynamic Jahn-Teller (DJT) [36] and Herzberg-Teller (HT) effects [37, 38] are incorporated into the VOF calculations.
@@
 We first describe the various steps of the computational protocol, and then we compare fluorescence lifetimes with measured values.
-Figure 1: Many-body states and optical spin polarization cycle of the NV- center in diamond.
+Figure 1: Many-body states and optical spin-polarization cycle of the NV- center in diamond.
 The spin-orbit splitting $\lambda_{z}$ represents the energy separation between the sublevels of the ${}^{3}\!E$ state.
 The intersystem crossing (ISC) rates from vibronic levels $\widetilde{A}_{1}$ and $\widetilde{E}_{1,2}$ to ${}^{1}\!A_{1}$ are labeled as $\Gamma_{A_{1}}$ (red dashed line) and $\Gamma_{E_{1,2}}$ (yellow dashed line).
-Radiative transitions ${}^{3}\!E\to{}^{3}\!A_{2}$ (blue solid line) and ${}^{1}\!A_{1}\to{}^{1}\!E$ (marron solid line), non-radiative transition ${}^{1}\!A_{1}\to{}^{1}\!E$ (marron dashed line), and ISC ${}^{1}\!E\to{}^{3}\!A_{2}$ (purple dashed line) are shown for completeness.
+Radiative transitions ${}^{3}\!E\to{}^{3}\!A_{2}$ (blue solid line) and ${}^{1}\!A_{1}\to{}^{1}\!E$ (maroon solid line), nonradiative transition ${}^{1}\!A_{1}\to{}^{1}\!E$ (maroon dashed line), and ISC ${}^{1}\!E\to{}^{3}\!A_{2}$ (purple dashed line) are shown for completeness.
 Theoretical framework—At temperatures near 0 K, the ISC rate $\Gamma_{A_{1}}$ is calculated using Fermi’s golden rule:
@@
 A detailed derivation of Eq.
-1 is provided in Sec.
-S5 [39].
+(1) is provided in Sec.
+S5 of Supplemental Material (SM) [39].
 The ISC rate $\Gamma_{E_{1,2}}$ can be similarly obtained by using the VOF for the $\widetilde{E}_{1,2}$ vibronic level [$F_{E_{1,2}}(\Delta)$].
@@
 The one-body SOC operators entering Eq.
-2 are defined as
+(2) are defined as
 $\hat{h}_{\mathrm{SO}}=\hat{\mathbf{z}}\cdot\hat{\mathbf{s}}=\sum_{I=1}^{N_{\mathrm{atom}}}\left(\hat{V}_{\mathrm{NL},I}^{\mathrm{FR}}-\hat{V}_{\mathrm{NL},I}^{\mathrm{SR}}\right),$ (3)
-where $\hat{V}_{\mathrm{NL},I}^{\mathrm{FR}}$ and $\hat{V}_{\mathrm{NL},I}^{\mathrm{SR}}$ are the non-local (NL) components of the fully relativistic (FR) and the scalar relativistic (SR) pseudopotentials, respectively, for atom $I$ in the supercell [42, 43, 44].
+where $\hat{V}_{\mathrm{NL},I}^{\mathrm{FR}}$ and $\hat{V}_{\mathrm{NL},I}^{\mathrm{SR}}$ are the nonlocal (NL) components of the fully relativistic (FR) and the scalar relativistic (SR) pseudopotentials, respectively, for atom $I$ in the supercell [42, 43, 44].
 Finite-size effects are addressed by converging the values of $\lambda_{z/\perp}$ with supercell size.
@@
 First, contributions from $a_{1}$-type phonon modes, $F^{a_{1}}(\Delta)$, are calculated using the Huang-Rhys (HR) theory [50, 20, 21].
-This calculation uses the atomic geometries of ${}^{3}\!E$ and ${}^{1}\!A_{1}$ and the phonon modes of ${}^{1}\!A_{1}$, derived from spin-conserving and spin-flip TDDFT calculations [22], which effectively include the multi-configurational nature of these states.
+This calculation uses the atomic geometries of ${}^{3}\!E$ and ${}^{1}\!A_{1}$ and the phonon modes of ${}^{1}\!A_{1}$, derived from spin-conserving and spin-flip TDDFT calculations [22], which effectively include the multiconfigurational nature of these states.
 Next, the DJT effect of ${}^{3}\!E$, involving $e$-type phonon modes, is incorporated to obtain the final $F_{A_{1}/E_{1,2}}(\Delta)$ [17].
 The computed VOFs are extrapolated to the dilute limit using the force constant embedding approach [20, 27].
-Equation 1 assumes the validity of the Franck-Condon (FC) principle and treats $\lambda_{\perp}$ as independent of atomic vibrations.
+Equation (1) assumes the validity of the Franck-Condon (FC) principle and treats $\lambda_{\perp}$ as independent of atomic vibrations.
 Here, we remove this assumption by accounting for the HT effect [37, 38] and compute the derivative of $\lambda_{\perp}$ with respect to configuration coordinates in the calculation of $F^{a_{1}}(\Delta)$.
@@
 Further details on our computational framework are in Secs.
-S1 and S2 [39].
-Results—We first compute the SOC parameter, $\lambda_{z/\perp}$, using many-body wavefunctions derived from group theory, using the minimal basis set of defect orbitals $a_{1}$, $e_{x}$, and $e_{y}$ (see Tables S1 and S2 [39]).
-The resulting expressions are: $\lambda_{z}=|\langle e_{x}|\hat{\mathbf{z}}|e_{y}\rangle|/2$ and $\lambda_{\perp}=|\langle a_{1}|\hat{\mathbf{z}}|e_{x}\rangle|/2$.
+S1 and S2 of SM [39].
+Although not explored here, this computational framework is readily applicable to ISC transitions from ${}^{1}\!E$ to ${}^{3}\!A_{2}$ ($t_{\pm}$ and $t_{z}$ in Fig.
+1) by substituting the SOC matrix element, the VOF, and the energy gap with those between ${}^{1}\!E$ and ${}^{3}\!A_{2}$.
+Results—We first compute the SOC parameter, $\lambda_{z/\perp}$, using many-body wave functions derived from group theory, using the minimal basis set of defect orbitals $a_{1}$, $e_{x}$, and $e_{y}$ (see Tables S1 and S2 in SM [39]).
+The resulting expressions are – $\lambda_{z}=|\langle e_{x}|\hat{\mathbf{z}}|e_{y}\rangle|/2$ and $\lambda_{\perp}=|\langle a_{1}|\hat{\mathbf{z}}|e_{x}\rangle|/2$.
 Results obtained from these expressions are denoted as $\lambda_{z/\perp}^{\mathrm{DFT}}$.
@@
 While QDET wave functions resemble those from the group theory analysis, they do contain additional configurations that introduce small yet important corrections, leading to a notable decrease in $\lambda_{z/\perp}^{\mathrm{QDET}}$ compared to $\lambda_{z/\perp}^{\mathrm{DFT}}$ (details in Secs.
-S3 and S4 [39]).
+S3 and S4 [39] of SM).
 By accounting for the weak dependence of $\lambda_{z/\perp}^{\mathrm{QDET}}$ on the active space choice and applying a finite-size correction (assumed to be the same as for $\lambda_{z/\perp}^{\mathrm{DFT}}$), we determine the final SOC parameters as $\lambda_{z,0}^{\mathrm{QDET}}=23.1\pm 4.3$ GHz and $\lambda_{\perp,0}^{\mathrm{QDET}}=29.4\pm 2.8$ GHz.
@@
 Using $p=0.304$ from Ref.
-[17], we obtain $p\lambda_{z,0}^{\mathrm{QDET}}=7.0\pm 1.3$ GHz, in good agreement with the measured values: $5.33\pm 0.03$ GHz [46, 47], whereas $p\lambda_{z,0}^{\mathrm{DFT}}$ overestimates experiment by a factor of two.
-Our computed $\lambda_{\perp,0}^{\mathrm{QDET}}$ also aligns well with the experimental estimate based on $\lambda_{\perp}/\lambda_{z}=1.2\pm 0.2$ relation [48, 49], while $\lambda_{\perp,0}^{\mathrm{DFT}}$ is approximately three times as large.
-We note that previous quantum chemistry CASSCF calculations [29] significantly underestimated $\lambda_{\perp}$, likely due to an inaccurate description of the NV- center’s many-body wavefunctions with a small C33H36N- cluster.
+[17], we obtain $p\lambda_{z,0}^{\mathrm{QDET}}=7.0\pm 1.3$ GHz, in good agreement with the measured values, $5.33\pm 0.03$ GHz [46, 47], whereas $p\lambda_{z,0}^{\mathrm{DFT}}$ overestimates experiment by a factor of 2.
+Our computed $\lambda_{\perp,0}^{\mathrm{QDET}}$ also aligns well with the experimental estimate based on $\lambda_{\perp}/\lambda_{z}=1.2\pm 0.2$ relation [48, 49], while $\lambda_{\perp,0}^{\mathrm{DFT}}$ is approximately 3 times as large.
+We note that previous complete active space self-consistent field (CASSCF) calculations [29] significantly underestimated $\lambda_{\perp}$, likely due to an inaccurate description of the NV- center’s many-body wave functions with a small C33H36N- cluster.
 As mentioned above, in our calculations of VOFs for the $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ ISC transitions, we first considered contributions from $a_{1}$-type phonon modes.
@@
 Here, we compute $F^{a_{1}}(\Delta)$ using the HR theory [50], based on equilibrium atomic geometries of the ${}^{3}\!E$ and ${}^{1}\!A_{1}$ states and $a_{1}$-type phonon modes of ${}^{1}\!A_{1}$, obtained from spin-conserving and spin-flip TDDFT calculations.
-A scaling factor is applied to adjust the HR factors, correcting for underestimated atomic displacements due to the use of the PBE functional [27, 22] (see Sec.
-S6 [39]).
+A scaling factor is applied to adjust the HR factors, correcting for underestimated atomic displacements due to the use of the Perdew-Burke-Emzerhof (PBE) functional [27, 22] (see Sec.
+S6 of SM [39]).
 We find notable differences between $F^{a_{1}}(\Delta)$ for ${}^{3}\!E\to{}^{1}\!A_{1}$ and ${}^{3}\!E\to{}^{3}\!A_{2}$ transitions [Fig.
@@
 However, we find that $\lambda_{\perp}$ varies by 8% between the ${}^{1}\!A_{1}$ and ${}^{3}E$ equilibrium atomic geometries (see Sec.
-S7 [39]).
+S7 of SM [39]).
 This variation is accounted for by including the HT effect [37, 38] in our calculations.
@@
 The DJT effect mixes the $A_{1}$, $A_{2}$, and $E_{1,2}$ electronic levels of ${}^{3}\!E$, giving rise to the vibronic states $\widetilde{A}_{1}$, $\widetilde{A}_{2}$, and $\widetilde{E}_{1,2}$ through coupling with $e$-type phonons.
-This coupling reduces the intensity of $F_{A_{1}}(\Delta)$ relative to the bare $F^{a_{1}}(\Delta)$ and generates a non-zero $F_{E_{1,2}}(\Delta)$, thereby enabling the $\Gamma_{E_{1,2}}$ ISC transition.
+This coupling reduces the intensity of $F_{A_{1}}(\Delta)$ relative to the bare $F^{a_{1}}(\Delta)$ and generates a nonzero $F_{E_{1,2}}(\Delta)$, thereby enabling the $\Gamma_{E_{1,2}}$ ISC transition.
 The onset of $F_{E_{1,2}}(\Delta)$ at 78 meV confirm its $e$-type phonon-assisted nature, consistent with the model of Ref.
@@
 Further details are provided in Sec.
-S5 [39].
+S5 of SM [39].
 We now combine the calculations of the SOC parameter $\lambda_{\perp,0}^{\mathrm{QDET}}$ and the VOFs $F_{A_{1}}(\Delta)$ and $F_{E_{1,2}}(\Delta)$ to compute the ISC rates $\Gamma_{A_{1}}$ and $\Gamma_{E_{1,2}}$ as functions of the energy gap $\Delta$.
@@
 4(a), while those for $\Gamma_{E_{1,2}}$ are in Fig.
-S10 [39].
+S10 of SM [39].
 Comparing the computed $\Gamma_{A_{1}}$ with experimental data [52, 48] yields an estimated $\Delta$ range of $[0.334,0.389]$ eV.
@@
 The inverse dependence of $\Gamma_{A_{1}/E_{1,2}}$ on $\Delta$ suggests that the NV- center’s ISC rates, and hence the optical spin-readout contrast, could be enhanced by decreasing $\Delta$, potentially through strain engineering [53, 54].
-Our QDET calculations predict a $\Delta$ value of 0.414 eV, which exceeds the upper bound of $\Delta$ shown in Fig.
-4(a), resulting in a $\Gamma_{A_{1}}$ rate of $66\pm 13$ MHz.
-This rate is slightly lower than the measured value of $100.5\pm 3.8$ MHz [48].
+Our QDET calculations predict a $\Delta$ value of 0.293 eV, which underestimates the lower bound of $\Delta$ shown in Fig.
+4(a), resulting in a $\Gamma_{A_{1}}$ rate of $180\pm 34$ MHz.
+This rate is higher than the measured value of $100.5\pm 3.8$ MHz [48].
 We also compare previously reported $\Delta$ values from various theoretical approaches [35, 55, 56, 57, 28, 29, 58] and compute the corresponding $\Gamma_{A_{1}}$ rates using $\lambda_{\perp,0}^{\textrm{QDET}}$ and $F_{A_{1}}(\Delta)$ from this work, as shown in Fig.
 4(b).
-Notably, different theoretical approaches predict $\Delta$ in a broad range of $[0.04,0.51]$ eV, leading to $\Gamma_{A_{1}}$ between 20 MHz and 450 MHz.
+Notably, different theoretical approaches predict $\Delta$ in a broad range of $[0.04,0.51]$ eV, leading to $\Gamma_{A_{1}}$ between 20 and 450 MHz.
 The values of $\Delta$ obtained from quantum chemistry approaches using carbon clusters are sensitive to both the level of theory and cluster size, resulting in $\Gamma_{A_{1}}$ variations by an order of magnitude.
@@
 (See Sec.
-S9 [39] for a detailed comparison of ISC rates evaluated here and in previous studies [48, 49, 17, 29]).
+S9 of SM [39] for a detailed comparison of ISC rates evaluated here and in previous studies [48, 49, 17, 29]).
 Figure 4: Intersystem crossing (ISC) rates and fluorescence lifetimes.
@@
 A visualization width of 0.02 eV is applied to each $\Delta$.
-Detailed data are provided in Table S5 [39].
+Detailed data are provided in Table S5 of SM [39].
 (c) Computed fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ compared with experimental values from this work and previous studies by Batalov et al.
@@
 Finally, we extend our computational framework for ISC rate calculations to finite temperatures ($T$) by summing contributions from different vibronic levels of ${}^{3}\!E$, weighted by the Boltzmann factor (see Sec.
-S8 [39]).
-We compute the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 K to 700 K, defined as $\tau(T)=1/\left[\Gamma_{\mathrm{Rad}}+\Gamma_{\mathrm{ISC}}(T)\right]$, where $\Gamma_{\mathrm{Rad}}=82.9$ MHz is the radiative decay rate of ${}^{3}\!E$ [52], and $\Gamma_{\mathrm{ISC}}(T)$ is the average ISC rate at a given $T$.
-To validate our calculations, we measured the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 K to 300 K in 50 K intervals.
+S8 of SM [39]).
+This approach effectively accounts for the phonon-induced orbital averaging effect and yields the averaged ISC rate from the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ [48].
+We compute the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 to 700 K, defined as $\tau(T)=1/\left[\Gamma_{\mathrm{Rad}}+\Gamma_{\mathrm{ISC}}(T)\right]$, where $\Gamma_{\mathrm{Rad}}=82.9$ MHz is the radiative decay rate of ${}^{3}\!E$ [52], and $\Gamma_{\mathrm{ISC}}(T)$ is the average ISC rate at a given $T$.
+To validate our calculations, we measured the fluorescence lifetimes of the $|m_{S}|=1$ sublevels of ${}^{3}\!E$ from 100 to 300 K in 50 K intervals.
 Combining these with previous measurements from Batalov et al.
@@
 Details on the experimental setup and lifetime fitting are provided in Secs.
-S10 and S11 [39].
-Despite moderate uncertainties in measurements and lifetime-fitting, our computed lifetimes agree well with experimental values below 600 K, as shown in Fig.
+S10 and S11 of SM [39].
+Despite moderate uncertainties in measurements and lifetime fitting, our computed lifetimes agree well with experimental values below 600 K, as shown in Fig.
 4(c).
@@
 Additionally, our framework can be adapted to calculate phosphorescence rates for transition metal impurities in solids [65], further expanding the applicability of first-principles calculations in predicting spin-related phenomena in quantum materials.
-Acknowledgments—The authors would like to thank Yuhang Ai for helpful discussions.
+Acknowledgments—The authors would like to thank Siyuan Chen, Victor Wen-zhe Yu, and Yuhang Ai for helpful discussions.
 The theoretical and computational work was primarily supported by the Midwest Integrated Center for Computational Materials (MICCoM) as part of the Computational Materials Sciences Program funded by the U.S.
-Department of Energy (Y.J., J.P., C.E.
-B.H., M.G., F.J.H., and G.G.).
+Department of Energy (Y.J., J.P., C.E., B.H., M.G., F.J.H., and G.G.).
 J.P.
@@
 Department of Energy, Office of Science, Basic Energy Sciences Materials Sciences and Engineering Division (B.P.
-and D.D.A.), the SNU-Global Excellence Research Center establishment project and the NRF grant funded by the Korean government (D.D.O.) (MSIT) No.
+and D.D.A.), the SNU-Global Excellence Research Center establishment project, and the NRF grant funded by the Korean government (D.D.O.) (MSIT) No.
 RS-2023-00258359, and from the AFOSR MURI under award No.
@@
 ALCC-ERCAP0025950, and resources of the University of Chicago Research Computing Center.
+Data availability—The data that support the findings of this Letter are openly available [66].